OpenStudy (anonymous):

The sets of C and D are defined as follows. C= {v| v<3} D= {v|v greater than or equal to 5} Write ∪CD and ∩CD using interval notation. If the set is empty, write ∅

OpenStudy (zzr0ck3r):

The union will be all the collection of all the elements that are in either set. \((-\infty, 3)\cup[5,\infty)\) The intersection will be all the elements that are in both sets, so all the elements the sets share with each other. \((-\infty, 3)\cap[5,\infty)=\{\}=\emptyset\)

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (zzr0ck3r):

np